### v1

@@ A.2 UK regions
We gathered data for the year 2019 for the 374 local authority districts (LADs) in the United Kingdom. Most variables were sourced from the Office of National Statistics (ONS) using API access via the R package onsr (Vasilopoulos, 2023). Some series are not directly accessible via API and were downloaded from the official websites of the ONS, Northern Ireland Statistics and Research Agency (NISRA) and the Scottish and Welsh government websites.
@@ A.3 EU regions
All the EU series for the year 2019 are collected from eurostat, an official database of the the European Union. We access the series at the NUTS2 and NUTS3 level via API calls using the R package eurostat (Lahti et al., 2017; Lahti et al., 2023). Note that fewer series are available on the NUTS3 level.
@@ A.5 US listed firms
We collected financial information as of end 2022131313We could not access data from 2019, as we do for the other datasets, because the free API access is limited to the most recent years., including balance sheet, income and share price data, of US listed firms using yahoo!finance’s [^14] https://finance.yahoo.com/ free API access. We excluded values on the variables return on revenue and return on equity when firms have negative values on stockholders’ equity or total revenue, respectively.
@@ Appendix B Methodology
B.1 Prompting
Additionally to the baseline completion prompting strategy (see Section 2.2), we also tested three alternative strategies. First, we apply the standard prompting template that is provided via the apply_chat_template function in Transformers package (Wolf et al., 2020), referred as question-answer prompting in this paper. For example, What was the unemployment rate in Orange County, California in 2019? This led to a lower response rate for US listed firms, with the model occasionally refusing to give an answer. Second, we test 5-shot prompting, where we provide 5 randomly chosen Q&A pairs answers to the model to learn from. Finally, we use chain-of-thought prompting, adding phrase ”think carefully before giving an answer” to the system prompt. The system prompt of the different prompting strategies are shown in Table AI.
Table LABEL:tab:response_rate indicates the response rates and figure AI compares the performance of different prompting strategies. On average, the completion strategy has the highest response rate and performs best.
@@ B.1 Prompting
Figure AI: Performance of LME when using different prompting strategies.
B.2 Extraction of numeric estimates from text output
Extracting numeric values from text outputs of smaller, open-source LLMs is challenging due to inconsistencies in the response structure and the numerical representation. We develop a rich extraction function with regular expressions. Firstly, we remove date strings (as these also contain numbers) and format numeric values by removing digit group separators before we convert answers into consistent units for the different variables. Additionally, with Q&A prompting (see XXX), LLMs typically provide the answer first and then offer an explanation. In contrast, for other strategies, LLMs tend to explain first and return the answer at the end. Therefore, based on the prompting strategy, we apply an additional conditional statement to select either the first or the last numeric value as the final prediction.
@@ Appendix C Additional results
Table AIII: Cross-validation performance of baseline model. The best performing approach in each row is highlighted in bold.
Variable Transfor No PCA PCA components Text
mation 5 25 100 200
@@
life expectancy at birth no 0.73 0.32 0.1 -0.28 -0.43 0.85
@@
population log 0.77 0.76 0.59 0.34 -0.21 0.72
risk-of-poverty rate no 0.60 0.29 0.01 -0.19 N/A 0.51
unemployment rate no 0.81 0.74 0.11 -0.08 -0.16 0.88
number of burglaries per hundred thousand inhabitants no 0.62 0.55 0.05 0.07 -0.23 -0.43
EU (NUTS3)
[…]
population log 0.82 0.85 0.78 0.68 0.55 0.65
number of burglaries per hundred thousand inhabitants no 0.52 0.24 0.14 0.17 0.27 -0.27
German districts
average age of the population no 0.82 0.71 0.81 0.72 0.51 0.61
[…]
population density (per square km) log 0.90 0.87 0.87 0.81 0.62 0.83
[…]
unemployment rate (%) no 0.83 0.80 0.8 0.78 0.54 0.75
number of hospital beds per 1,000 inhabitants no 0.54 0.45 0.50 0.48 0.31 N/A
number of business registrations per 10,000 inhabitants no 0.65 0.52 0.64 0.56 0.35 0.29
number of government employees per 1,000 inhabitants no 0.64 0.43 0.61 0.59 0.46 0.11
number of completed housing units per 1,000 inhabitants no 0.47 0.33 0.43 0.31 0.28 -0.10
number of filed corporate insolvencies per 10,000 taxable companies no 0.48 0.45 0.46 0.40 0.35 0.32
percentage change in population compared to the previous year no 0.59 0.57 0.52 0.44 0.39 0.48
percentage of school leavers with a general university entrance qualification (%) no 0.65 0.55 0.65 0.52 0.33 0.07
UK districts
average house price (in £) no 0.96 0.87 0.73 0.09 0.03 0.94
GDP per capita (in £) log 0.62 0.42 0.54 0.48 0.21 0.48
life expectancy for females at birth no 0.80 0.66 0.75 0.66 0.38 0.70
median age no 0.90 0.83 0.85 0.63 0.48 0.85
median annual gross pay (in £) no 0.78 0.77 0.66 0.54 -0.03 0.63
population log 0.86 0.79 0.77 0.65 0.37 0.89
unemployment rate no 0.75 0.61 0.61 0.39 0.05 0.7
percentage of the population that live in income deprivation no 0.87 0.84 0.85 0.76 0.44 0.85
US counties
GDP per capita (in \$) log 0.61 0.46 0.54 0.43 0.42 0.46
life expectancy at birth no 0.75 0.75 0.71 0.64 0.55 0.71
median monthly two-bedroom rent (in \$) log 0.75 0.70 0.71 0.52 0.50 0.60
personal income per capita (in \$) no 0.73 0.66 0.68 0.54 0.40 0.61
population log 0.87 0.81 0.82 0.83 0.82 0.90
unemployment rate no 0.63 0.49 0.48 0.37 0.19 0.65
number of business establishments per 100,000 people log 0.68 0.61 0.66 0.66 0.63 0.03
percentage change in population compared to the previous year no 0.52 0.44 0.45 0.38 0.27 0.41
proportion of mortgages being 90 or more days delinquent no 0.71 0.65 0.65 0.51 0.34 0.41
US firms
cost-to-revenue ratio log 0.61 0.07 0.46 0.56 0.55 -0.03
debt-to-equity ratio cubic 0.37 0.36 0.32 0.29 0.23 0.32
market capitalization log 0.91 0.89 0.90 0.91 0.90 0.83
price-earnings ratio cubic 0.5 0.46 0.49 0.46 0.45 0.37
price-to-book ratio cubic 0.28 0.14 0.25 0.19 0.11 0.39
return on assets cubic 0.69 0.64 0.68 0.68 0.68 0.45
return on equity cubic 0.67 0.62 0.66 0.65 0.65 0.44
return on revenue cubic 0.61 0.60 0.59 0.58 0.57 0.48
total assets log 0.93 0.88 0.92 0.93 0.93 0.81
@@
Figure AIII: Performance of base model in US firms dataset as a function of number of by PCA components.
Figure AIV: Performance of base model in US firms dataset as a function of number of by PCA components.
@@
Figure AVIII: Super-resolution results for Phi3 3.8B and Llama 3.2. 1B.

### v2

@@ A.2 UK regions
We gathered data for the year 2019 for the 374 local authority districts (LADs) in the United Kingdom. Most variables were sourced from the Office for National Statistics (ONS) using API access via the R package onsr (onsr). Some series are not directly accessible via API and were downloaded from the official websites of the ONS, Northern Ireland Statistics and Research Agency (NISRA) and the Scottish and Welsh government websites.
@@ A.3 EU regions
All the EU series for the year 2019 are collected from Eurostat, an official database of the European Union. We access the series at the NUTS2 and NUTS3 levels via API calls using the R package eurostat (lahti2017; eurostat). Note that fewer series are available on the NUTS3 level.
@@ A.5 US listed firms
We collected financial information as of the end of 2022161616We could not access data from 2019, as we do for the other datasets, because the free API access is limited to the most recent years., including balance sheet, income and share price data, of US listed firms using the Yahoo Finance [^17] https://finance.yahoo.com/ free API access. We excluded values on the variables return on revenue and return on equity when firms have negative values on stockholders’ equity or total revenue, respectively.
@@ Appendix B Methodology
B.1 Computational requirements
We used an NVIDIA Tesla V100 with 16 GB of video memory to generate the LLM’s embeddings and generate text. With 16-bit (fp16) parameters, this memory was sufficient for all models except Llama 70B variant. Using low-priority virtual machines, compute costs were under US \$0.80 per hour for the GPU. Our baseline Llama 3 8B produced embeddings for a query in  0.1 seconds. We also tested the speed of a quantised version of Llama 3 8B running locally on an AMD Ryzen 7 3700X CPU. This required  0.9 seconds per query.
B.2 Prompting
In addition to the baseline completion prompting strategy (see Section 2.2), we also tested three alternative strategies. First, we apply the standard prompting template that is provided via the apply_chat_template function in the Transformers package (wolf2020transformers), referred to as question-answer prompting in this paper. For example, What was the unemployment rate in Orange County, California in 2019? This led to a lower response rate for US listed firms, with the model occasionally refusing to give an answer. Second, we test 5-shot prompting, where we provide 5 randomly chosen Q&A pairs to the model to learn from. Finally, we use chain-of-thought prompting, adding the phrase “think carefully before giving an answer” to the system prompt. The system prompts for the different prompting strategies are shown in Table AI.
Table LABEL:tab:response_rate indicates the response rates and Figure AI compares the text output’s performance of different prompting strategies. On average, the completion strategy has the highest response rate and performs best.
@@ B.2 Prompting
Figure AI: Performance of text output when using different prompting strategies.
B.3 Extraction of numeric estimates from text output
Extracting numeric values from text outputs of smaller, open-source LLMs is challenging due to inconsistencies in the response structure and the numerical representation. We develop a rich extraction function with regular expressions. Firstly, we remove date strings (as these also contain numbers) and format numeric values by removing digit group separators before we convert answers into consistent units for the different variables. Additionally, with Q&A prompting, LLMs typically provide the answer first and then offer an explanation. In contrast, for other strategies, LLMs tend to explain first and return the answer at the end. Therefore, based on the prompting strategy, we apply an additional conditional statement to select either the first or the last numeric value as the final prediction.
@@ Appendix C Additional results
Table AIII: Cross-validation performance of the baseline LLM comparing LME with text output. For each performance metric, the best-performing approach is shown in bold. LME performance is based on 25 iterations of 5-fold cross-validation. Numbers before the parentheses indicate the mean; numbers in parentheses show the minimum and maximum performance across the 25 iterations.
Spearman Pearson
Variable Transfor- LME Text LME Text
mation
US firms
return on assets cubic 0.69 (0.69-0.70) 0.45 0.76 (0.75-0.76) 0.55
return on revenue cubic 0.60 (0.60-0.61) 0.48 0.66 (0.66-0.67) 0.54
market capitalization log 0.91 (0.91-0.91) 0.83 0.91 (0.91-0.91) 0.50
return on equity cubic 0.67 (0.66-0.67) 0.44 0.68 (0.68-0.68) 0.50
total assets log 0.94 (0.93-0.94) 0.81 0.92 (0.92-0.92) 0.48
price-earnings ratio cubic 0.50 (0.49-0.51) 0.37 0.46 (0.45-0.47) 0.21
price-to-book ratio cubic 0.28 (0.26-0.29) 0.39 0.14 (0.11-0.15) 0.12
debt-to-equity ratio cubic 0.37 (0.36-0.38) 0.32 0.18 (0.15-0.19) 0.08
cost-to-revenue ratio log 0.60 (0.59-0.62) -0.03 0.48 (0.46-0.49) 0.02
US counties
population log 0.87 (0.86-0.87) 0.90 0.89 (0.88-0.89) 0.91
life expectancy at birth no 0.75 (0.74-0.76) 0.71 0.74 (0.73-0.74) 0.69
unemployment rate no 0.63 (0.60-0.67) 0.65 0.63 (0.60-0.67) 0.64
median monthly two-bedroom rent (in \$) log 0.75 (0.74-0.75) 0.60 0.85 (0.84-0.85) 0.63
personal income per capita (in \$) no 0.73 (0.70-0.73) 0.61 0.72 (0.71-0.73) 0.63
GDP per capita (in \$) log 0.60 (0.59-0.62) 0.46 0.56 (0.53-0.57) 0.42
proportion of mortgages being 90 or more days delinquent no 0.70 (0.66-0.72) 0.41 0.69 (0.65-0.71) 0.41
percentage change in population compared to the previous year no 0.51 (0.50-0.52) 0.41 0.48 (0.47-0.49) 0.38
number of business establishments per 100,000 people log 0.69 (0.68-0.70) 0.03 0.68 (0.67-0.69) 0.04
UK districts
average house price (in £) no 0.94 (0.91-0.96) 0.94 0.89 (0.86-0.92) 0.93
population log 0.85 (0.84-0.86) 0.89 0.85 (0.81-0.86) 0.91
median age no 0.89 (0.88-0.91) 0.85 0.89 (0.88-0.91) 0.82
percentage of the population that live in income deprivation no 0.87 (0.85-0.88) 0.85 0.86 (0.84-0.88) 0.81
median annual gross pay (in £) no 0.77 (0.76-0.80) 0.63 0.83 (0.81-0.84) 0.72
life expectancy for females at birth no 0.80 (0.76-0.82) 0.70 0.76 (0.70-0.81) 0.70
unemployment rate no 0.75 (0.70-0.78) 0.70 0.73 (0.67-0.76) 0.68
GDP per capita (in £) log 0.63 (0.59-0.69) 0.48 0.64 (0.58-0.69) 0.66
German districts
number of hospital beds per 1,000 inhabitants no 0.53 (0.51-0.55) 0.52 (0.49-0.55) 0.00181818With constant responses across districts, the sample correlation is not defined. We record performance as 0.
population density (per square km) log 0.90 (0.88-0.90) 0.83 0.91 (0.90-0.92) 0.82
unemployment rate (%) no 0.83 (0.82-0.84) 0.75 0.81 (0.79-0.83) 0.74
[…]
average age of the population no 0.82 (0.81-0.84) 0.61 0.84 (0.83-0.85) 0.57
[…]
percentage change in population compared to the previous year no 0.57 (0.54-0.60) 0.48 0.57 (0.52-0.61) 0.42
number of business registrations per 10,000 inhabitants no 0.63 (0.57-0.66) 0.29 0.66 (0.59-0.69) 0.36
number of filed corporate insolvencies per 10,000 taxable companies no 0.49 (0.45-0.52) 0.32 0.59 (0.56-0.62) 0.34
percentage of school leavers with a general university entrance qualification (%) no 0.65 (0.63-0.67) 0.07 0.63 (0.60-0.65) 0.13
number of government employees per 1,000 inhabitants no 0.63 (0.62-0.65) 0.11 0.63 (0.61-0.65) 0.08
number of completed housing units per 1,000 inhabitants no 0.47 (0.41-0.52) -0.10 0.42 (0.35-0.46) 0.04
[…]
@@
population log 0.73 (0.69-0.77) 0.72 0.75 (0.71-0.81) 0.68
risk-of-poverty rate no 0.55 (0.35-0.61) 0.51 0.47 (0.34-0.59) 0.50
@@
number of burglaries per hundred thousand inhabitants no 0.55 (0.27-0.62) -0.43 0.47 (0.28-0.56) -0.01
@@
Refer to captionRefer to caption
Figure AIII: Performance of the base model on the US counties dataset as a function of the number of PCA components. LME’s mean performance across 25 cross-validation iterations is shown as a line; the shaded band indicates the minimum–maximum range across iterations.
Figure AIV: Performance of the base model on the US-firms dataset as a function of the number of PCA components. LME’s mean performance across 25 cross-validation iterations is shown as a line; the shaded band indicates the minimum–maximum range across iterations.